# Human limits in Machine Learning: Prediction of plant phenotypes using soil microbiome data

### v1

@@ Abstract
The preservation of soil health has been identified as one of the main challenges of the XXI century given its vast (and potentially threatening) ramifications in agriculture, human health and biodiversity. Here, we provide the first deep investigation of the predictive potential of machine-learning models to understand the connections between soil and biological phenotypes. Indeed, we investigate an integrative framework performing accurate machine-learning-based prediction of plant phenotypes from biological, chemical and physical properties of the soil via two models: random forest and Bayesian neural network. We show that prediction is improved, as evidenced by higher weighted F1 scores, when incorporating into the models environmental features like soil physicochemical properties and microbial population density in addition to the microbiome information. Furthermore, by exploring multiple data preprocessing strategies such as normalization, zero replacement, and data augmentation, we confirm that human decisions have a huge impact on the predictive performance. In particular, we show that the naive total sum scaling normalization that is commonly used in microbiome research is not the optimal strategy to maximize predictive power. In addition, we find that accurately defined labels are more important than normalization, taxonomic level or model characteristics. That is, if humans are unable to classify the samples and provide accurate labels, the performance of machine-learning models will be limited. Lastly, we present strategies for domain scientists via a full model selection decision tree to identify the human choices that maximize the prediction power of the models. Our work is accompanied by open source reproducible scripts (https://github.com/solislemuslab/soil-microbiome-nn) for maximum outreach among the microbiome research community.
Appendix A Introduction
It is undeniable that machine learning (ML) has revolutionized the manner in which scientific research is performed in recent years. Among the many tasks performed by ML models in our daily lives, researchers have relied on ML to assist in clinical diagnoses [13], identification of bacterial phenotypes such as antimicrobial resistance [3], and even identification of objects in space [15]. Recently, the vast evidence of the prediction power of ML models on a wide range of applications has launched the adoption of these models in other domains such as sustainable agriculture where soil health – characterized by a wide range of biological, chemical, and physical properties [28] – is explored as the potential driver to predict plant and crop phenotypes, such as disease susceptibility or yield.
@@ Appendix A Introduction
Fortunately, the rapid development of ML models could fill this gap. Over the years, ML models have been used on high-dimensional, highly interconnected data for prediction and decision-making that does not rely on a priori knowledge on the effects and interactions among covariates. Since we have little or no knowledge of most species among the thousands contained in the soil microbial communities, it would be beneficial to exploit the power of ML models on microbiome data as they are able to explore the unknown interactions between the microbial communities and crop phenotypes.
@@
Here, we explore the potential of ML models in the prediction of crop phenotypes of yield and disease from soil data. We utilize a real dataset from potato fields in Wisconsin and Minnesota and focus on the performance of the models while facing data challenges related to binarization, imbalance, compositionality, sparsity and high dimensionality. Furthermore, we test the impact of specific data preprocessing steps such as 1) different normalization and zero replacement strategies to overcome the compositionality and sparsity, 2) different feature selection strategies to overcome the high dimensionality, and 3) data augmentation techniques to overcome the imbalance of the data. In addition, we are also interested in answering the question of whether soil microbiome data alone has enough predictive power to predict the disease and yield phenotypes or whether other information on the soil, such as chemical composition, is necessary for accurate prediction. The answer to this question will inform farmers on the best data collection strategies given that soil microbiome data is expensive. That is, if the soil microbiome data does not provide enough predictive power compared to other (cheaper) soil measurements, then the collection of microbiome data for prediction purposes would be futile.
For our investigation, we select two distinct ML models: 1) Random forests (RF) as they have always been identified as the top predicting models in a variety of domain applications [5]. 2) Neural networks given that they are capable of exploring the arbitrary complex interactions among features that we do not have prior knowledge of. However, with the very small sample size and large feature space of this study, the traditional back-propagation-based neural network models will likely have a large bias or overfit [12]. Thus, we select the Bayesian neural network (BNN) as our second model, which offers natural protection against overfitting [9].
@@
Among the main findings, we can highlight that microbiome data alone indeed has predictive power for disease outcomes, especially for pitted scab disease, but not to predict yield. We also find that the most powerful prediction is achieved by combination of environmental information and microbiome data. Among the data preprocessing strategies that we explored, we find that normalization and zero replacement strategies have a huge impact on the prediction power of the models, yet there are strong interaction effects with taxonomic levels, and thus, it is impossible to identify one strategy that outperforms the others. In terms of the model, we found no clear differences between the RF and the BNN, yet the latter is more computationally expensive and prohibited on some datasets with a large number of predictors, which prompts us to prefer RF for most settings. We conclude our investigation with a full model selection (FMS) decision tree [25, 26] that allows us to identify combinations of normalization, zero replacement, feature selection, and model that yield the highest prediction accuracy and that can serve to provide specific strategies for other researchers with similar datasets.
Appendix B Results
Figure 1 shows a graphical representation of our pipeline with three major steps: (1) data preparation, (2) feature selection, and (3) classification based on random forest (RF) and Bayesian neural network (BNN) models with various types of predictors, including microbiome data, environmental, and a combination of both. Every panel in this graphical abstract will be described in more details in the Methods section.
@@ Appendix B Results
Figure 1: Workflow of the analyses with three main steps: (1) Data Preparation, (2) Feature Selection, and (3) Classification. In (1) Data Preparation, we consider OTUs at different taxonomic levels and filter by sample size. In addition, we perform five normalization methods and four zero replacement methods. In (2) Feature selection, we rank OTUs based on the number of times they are selected as important features by machine learning (ML) criteria. Then, we select OTUs that display the greatest degree of difference on microbial networks reconstructed from samples of each class. Finally, we score OTUs based on whether they are selected by ML ($score=1$), by network comparison ($score=2$), both ($score=3$) or neither ($score=0$). In (3) Classification, the Venn diagram depicts the different types of predictors: microbiome (OTUs), environmental (Env), and the combination of both. random forest and Bayesian neural network classification models are fitted on the different input predictors.
@@
B.1 Overall performance of predictive models: Manual binarization causes inaccurate prediction of yield
@@
First, we identify the outcomes (disease or yield) that are accurately predicted across models (and are thus robust to prediction regardless of model choices), as well as the models that accurately predict across outcomes (and are thus the most powerful model alternatives). To do so, we aggregate the weighted F1 scores on data preprocessing choices such as normalization, zero replacement, and taxonomic levels for every model and every outcome. Figure S3 (random forest) and Figure S4 (BNN) show the boxplots with the weighted F1 scores of the 14 models described in Table 3. Columns correspond to the six responses: four diseases and two yield outcomes. For a given panel (model in row and response in column), the boxplot corresponds to the different weighted F1 scores for every combination of normalizations/zero replacement strategies as well as different taxonomic levels (Table 4). For example, the boxplots for the ALL-OTU model (first row) include weighted F1 scores of the model fit on 20 normalization/zero replacement strategies, and 5 taxonomic levels (100 different weighted F1 scores per outcome). The dashed line in each panel corresponds to the average weighted F1 score of the model when fit with all random datasets (see Section B.2.1). This line allows us to assess whether the real data has more predictive power than random data.
@@
While these plots do not allow us to distinguish differences by taxonomic level or normalization/zero replacement strategy (more on that in the next subsections), we can identify outcomes (columns) that can be more accurately predicted across models (rows). Additionally, we can identify models that are capable of accurately predicting more outcomes. It is readily evident, for example, that the yield outcomes cannot be accurately predicted by any model as all the weighted F1 scores fall consistently below dashed line. It is notable that the yield responses are those for which there is not a clear binarization strategy. Since we are artificially separating samples into the two classes (low and high yield) based on whether they are above or below the variety-specific median, samples on the boundary will in fact be very similar to each other, and thus, difficult to classify. Furthermore, the poor prediction of yield is not restricted to one data type (microbiome vs environmental) which also suggests that the prediction challenges arise from the binarization process rather than the model or set of predictors.
@@
Disease outcomes, on the contrary, display higher weighted F1 scores overall, and in particular, pitted scab displays weighted F1 scores that are consistently above the random prediction dashed line across different models. For the case of black scurf, even when the weighted F1 scores are consistently above dashed line, this is a deceiving result, as this disease outcome is highly imbalanced. This means that a naive model predicting all samples to belong to the majority class will have high prediction accuracy (see dashed line above 0.8 for random data). We investigate the prediction of black scurf more carefully with the augmented data that balances the proportion of both classes (Section B.2.4). Finally, given the poor performance on yield, we focus on the fine-grained description of results for the disease outcomes only for the remaining of the manuscript.
Regarding differences by model (rows), the most accurate prediction across all outcomes can be obtained with alpha diversity and soil chemistry data (Alpha+Soil) for the case of RF (Figure S3) and with OTUs identified as important by both ML and network comparison strategies (more on feature selection in Section B.2.3) along with soil chemistry data (OTU-S3+Soil) for the case of BNN models (Figure S4). Both of these results show that incorporating environmental information in conjunction with microbiome data can increase the predictive power for diseases outcomes.
B.2 Prediction of potato disease from microbiome data
B.2.1 More power to predict pitted scab disease in microbiome data compared to random data
For the pitted scab on the top row, the F1 scores for the original dataset (red dots) are on the upper right quadrant compared to points from random datasets (blue dots). This means that the F1 scores on the original data are higher than F1 scores on random data, and this, the original data has more predictive power than random data. However, for the superficial scab on the bottom row, many random datasets show better performance (higher F1 score values) than the real data, and thus, the real data does not contain much information to predict this disease. Other diseases as well as yield outcomes also underperform compared to random data as shown in Figure S5 in the Appendix. We highlight that pitted scab is the most severe case of the potato disease common scab, and possibly a stronger signal of the soil microbial community composition and corresponding disease-suppressiveness than other measures of the disease such as superficial scab.
@@ B.2.1 More power to predict pitted scab disease in microbiome data compared to random data
B.2.2 No normalization or zero replacement strategy proven Superior in terms of predictive power
@@
One of the goals of our study is to identify ideal data preprocessing steps that are guaranteed to maximize predictive power on the machine learning models. Our results, however, suggest that there is not a clear pattern of superiority among the normalization or zero replacement strategies, and that selection of the data preprocessing steps needs to be data specific (and taxonomic level specific).
@@
Figure 3 shows the weighted F1 scores for pitted scab for different combinations of normalization and zero replacement strategies (x-axis) for the two types of models (RF and BNN). These analyses include all OTUs under the five taxonomic levels (different colors). Similar plots for other diseases are presented in Figures 6 to S10 in the Appendix. As mentioned, we note that there is not an apparent pattern that points at any normalization or zero replacement strategy being superior to other alternatives across all taxonomic levels. In fact, there is considerable interaction between the normalization/zero replacement method and the taxonomic level. For example, for the RF model, the best result is achieved with Phylum level and cumulative sum scaling normalization with pseudo-zero replacement strategy (CSS+pseudo) or common sum scaling normalization without any zero replacement strategy (COM+none). For BNN, however, the best results are achieved with the common sum scaling normalization with the multiplicative zero replacement (COM+multRepl) for the Phylum level.
@@
For a given normalization/zero replacement strategy (x-axis), the fact that the points in the scatterplot are not overlapping shows that taxonomic levels have a impact on the predictive power of the model. There is, however, no apparent pattern on a given taxonomic level providing more predictive signal. Phylum level seems to produce higher weighted F1 scores for some normalization/zero replacement strategies, but not all.
@@
When we compare the range of weighted F1 scores across normalization and zero replacement strategies, we see that the effect of the strategy is not negligible. For example, for a Phylum level, the lowest weighted F1 score is around 0.75 for centered log-ratio normalization with pseudo-zero replacement strategy (clr+pseudo) to around 0.9 for cumulative sum scaling normalization with pseudo-zero replacement strategy (CSS+pseudo). This implies that for a given taxonomic level, the resulted weighted F1 score will be highly influenced by the normalization and zero replacement strategy. Traditionally, microbiome researchers use the total sum scaling normalization without any zero replacement strategy (TSS+none) on their data which has a range of 0.80 to 0.90 weighted F1 scores for the RF model (0.8 to 0.85 for the BNN model) depending on the taxonomic level.
@@
The strong interaction effects of taxonomic level, normalization and zero replacement strategy prevent us from making recommendations about the best data preprocessing practices that can be generalizable to other datasets. We conclude by suggesting data practitioners to try different normalization and zero replacement strategies rather than just one, but see Section B.2.5 for more recommendations.
@@
Figure 3: Weighted F1 scores (y-axis) for random forest and Bayesian neural network (Bayesian NN) models for the pitted scab disease (Scabpit) under the 20 normalization/zero replacement strategies (x-axis). The lack of pattern prevents us from making recommendations of optimal strategies for microbiome data. We can conclude, however, that taxonomic levels, normalization and zero replacement strategies have an effect on the prediction accuracy of the models as evidenced by the broad range displayed by the points.
@@
B.2.3 Effective preservation of predictive signal with different feature selection strategies
One of the standard steps in the machine learning pipeline is feature selection, especially for cases of high-dimensional data. We compare the ability to retain predictive signal of three feature selection strategies: standard importance score from machine learning methods, comparison of microbial network topologies and combination of both. More details on the feature selection strategies can be found in Methods.
@@ B.2.3 Effective preservation of predictive signal with different feature selection strategies
Figure 4 shows the weighted F1 scores for the two types of models (RF and BNN) on pitted scab disease (Scapbit) under different subsets of predictors: (1) all OTUs (ALL-OTU), (2) only OTUs that were identified as important by the ML strategy (OTU-S1), (3) only OTUs that were identified as important by the network comparison strategy (OTU-S2), (4) OTUs that were identified as important by both strategies (OTU-S3), or (5) OTUs that were not identified as important by neither strategy (OTU-S0). For fair comparison, we include the same number of predictors in OTU-S0 as in OTU-S3. Similar figures for other responses are shown in Figures S11 to S15 in the Appendix.
Again, we perceive strong interaction between taxonomic level and feature selection strategy. For the RF model, the highest weighted F1 score is achieved when including all OTUs (ALL-OTU) at Order level whereas for the BNN model, the highest weighted F1 score is achieved when including OTUs identified by the ML strategy (OTU-S1) at Genus level. RF models on all OTUs (ALL-OTU) have weighted F1 score above 0.8 in all taxonomic levels which suggest that this model could be a better alternative compared to BNN which is more computationally intensive. There are also smaller differences in RF models when comparing the performance on OTU-S3 (important OTUs) and ALL-OTU (all OTUs) which suggests that the feature selection strategy is sufficient to preserve the predictive signal in the data while reducing the number of predictors in the model. This is relevant for computationally intensive models such as BNN that do not allow the inclusion of all OTUs for certain taxonomic levels.
@@
B.2.4 Robust prediction in imbalanced datasets with data augmentation
@@
To illustrate this, Figure 5 depicts the weighted F1 scores on original and augmented datasets for all yield and disease outcomes (x-axis) and both models (RF and BNN). The range of each box plot depicts the weighted F1 scores for 20 normalized datasets at each taxonomic level. We observe that black scurf and pitted scab can be reliably predicted across taxonomic levels as their median weighted F1 scores for all taxonomic orders is around $0.8$ when the models are fitted on the original datasets. As mentioned before, however, black scurf is highly imbalanced, so the results on the original data are not reliable. Fortunately, the median weighted F1 scores on augmented data (which is perfectly balanced by design) increases for both diseases, such that they are around $0.9$ for all taxonomic levels. These results suggest that data augmentation, especially in cases of highly imbalanced data, is an appropriate strategy that increases the accuracy of the prediction. One has to be careful, however, in that augmented data can yield certain models prohibited. For example, the BNN model could not be fit on the augmented datasets for Order, Family nor Genus levels due to computational limitations.
Figure 5: Boxplots with weighted F1 scores by model (rows): random forest (RF) and Bayesian neural network (BNN) for original and augmented OTU predictors (columns). The x-axis represents the yield and disease outcomes. The range of each box plot depicts the weighted F1 scores for 20 normalized datasets at each taxonomic level.
B.2.5 Identifying general practices to predict potato disease from microbiome data using full model selection models
@@
With a full model selection (FMS) [25, 26, 24] strategy, however, we are able to identify the choices that yield the highest measure of performance. More details on the FMS models can be found in Methods. Figures 6 and 7 show the FMS decision trees for the RF and BNN models on pitted scab disease, respectively. A FMS decision tree shows the different data preprocessing steps that yield different weighted F1 scores, so that practitioners can select the options that result in the highest predictive power. Here, we have five taxonomic levels, 20 normalization+zero replacement strategies, and 2 data augmentation options: no data augmentation (Aug=0) and data augmentation (Aug=1). Thus, in total, we have 200 data preprocessing options (20 normalization strategies times 5 taxonomic levels times 2 data augmentation).
@@
To interpret a FMS decision tree, each node corresponds to a specific step in the data preprocessing pipeline, for example, whether to perform data augmentation or not. If the condition is true, we follow the branch to the left; if the condition is false, we follow the branch to the right. At the top of the decision tree, we have the root which represents the data preprocessing step that has the greatest effect on model accuracy. At the bottom of the decision tree, we have the leaves with the average weighted F1 score of the model fitted on the data that satisfy all conditions towards the root. The darker the red, the higher the weighted F1 score. Each node also displays the percentage of data preprocessing options included in the node. For example, in Figure 6, the root node covers 100% of the options with average weighted F1 scores 0.865. The condition at the root node ($\text{Aug}=0$) represents the case of "no data augmentation". Thus, "true" (left of the root) means "no data augmentation", and "false" (right of the root) means "data augmentation". For simplicity, we denote the 20 normalization/zero replacement strategies as $\text{NM}_{i}$ for $i=1,\dots,20$. See Section D.1.2 for a description on each normalization/zero replacement strategy.

### v2

@@ Abstract
The preservation of soil health is a critical challenge in the 21st century due to its significant impact on agriculture, human health, and biodiversity. We provide the first deep investigation of the predictive potential of machine learning models to understand the connections between soil and biological phenotypes. We investigate an integrative framework performing accurate machine learning-based prediction of plant phenotypes from biological, chemical, and physical properties of the soil via two models: random forest and Bayesian neural network. We show that prediction is improved when incorporating environmental features like soil physicochemical properties and microbial population density into the models, in addition to the microbiome information. Exploring various data preprocessing strategies confirms the significant impact of human decisions on predictive performance. We show that the naive total sum scaling normalization that is commonly used in microbiome research is not the optimal strategy to maximize predictive power. Also, we find that accurately defined labels are more important than normalization, taxonomic level or model characteristics. In cases where humans are unable to classify samples accurately, machine learning model performance is limited. Lastly, we provide domain scientists via a full model selection decision tree to identify the human choices that optimize model prediction power.
Introduction
It is undeniable that machine learning (ML) has revolutionized the manner in which scientific research is performed in recent years. Among the many tasks performed by ML models in our daily lives, researchers have relied on ML to assist in clinical diagnoses [13], identification of bacterial phenotypes such as antimicrobial resistance [3], and even identification of objects in space [15]. Recently, the vast evidence of the prediction power of ML models on a wide range of applications has launched the adoption of these models in other domains such as sustainable agriculture where soil health – characterized by a wide range of biological, chemical, and physical properties [28] – is explored as an important driver to predict plant phenotypes, such as disease susceptibility or yield.
@@ Introduction
Fortunately, the rapid development of ML models could fill this gap. Over the years, ML models have been used on high-dimensional, highly interconnected data for prediction and decision-making that does not rely on a priori knowledge on the effects and interactions among covariates. Since we have little or no knowledge of most species among the thousands contained in the soil microbial communities, it would be beneficial to exploit the power of ML models on microbiome data as they are able to explore the unknown interactions between the microbial communities and plant phenotypes.
@@
Here, we explore the potential of ML models in the prediction of plant phenotypes of yield and disease from soil data. We utilize a real dataset from potato fields in Wisconsin and Minnesota and focus on the performance of the models while facing data challenges related to binarization, imbalance, compositionality, sparsity and high dimensionality. Furthermore, we test the impact of specific data preprocessing steps such as 1) different normalization and zero replacement strategies to overcome the compositionality and sparsity, 2) different feature selection strategies to overcome the high dimensionality, and 3) data augmentation techniques to overcome the imbalance of the data. In addition, we are also interested in answering the question of whether soil microbiome data alone has enough predictive power to predict the disease and yield phenotypes or whether other information on the soil, such as chemical composition, is necessary for accurate prediction. The answer to this question will inform farmers on the best data collection strategies given that soil microbiome data is expensive. That is, if the soil microbiome data does not provide enough predictive power compared to other (cheaper) soil measurements, then the collection of microbiome data for prediction purposes would be futile. Figure 1 shows a graphical description of our pipeline.
For our investigation, we have chosen two distinct machine learning models. First, we employ Random Forests (RF), which have consistently demonstrated superior predictive performance across various domains [5]. Further details about this method can be found in Section ‘Random Forest Model’. Second, we utilize a Bayesian Neural Network (Bayesian NN), known for its inherent protection against overfitting [9] (See Section ‘Bayesian Neural Network Model’). Given their capacity to capture complex interactions among features, neural networks are valuable, especially when dealing with unknown relationships. However, considering the small sample size and large feature space in our study, traditional back-propagation-based neural network models may exhibit a substantial bias or overfitting [12], hence the need to explore the Bayesian version.
@@
Among the main findings, we can highlight that microbiome data alone indeed has predictive power for disease outcomes, especially for pitted scab disease, but not to predict yield. We also find that the most powerful prediction is achieved by combination of environmental information and microbiome data. Among the data preprocessing strategies that we explored, we find that normalization and zero replacement strategies have a huge impact on the prediction power of the models, yet there are strong interaction effects with taxonomic levels, and thus, it is impossible to identify one strategy that outperforms the others. In terms of the model, we found no clear differences between the RF and the Bayesian NN, yet the latter is more computationally expensive and prohibited on some datasets with a large number of predictors, which prompts us to prefer RF for most settings. We conclude our investigation with a full model selection (FMS) decision tree [25, 26] that allows us to identify combinations of normalization, zero replacement, feature selection, and model that yield the highest prediction accuracy and that can serve to provide specific strategies for other researchers with similar datasets.
@@
Figure 1: Workflow of the analyses with three main steps: (1) Data Preparation, (2) Feature Selection, and (3) Classification. In (1) Data Preparation, we consider OTUs (number of OTUs = N) at different taxonomic levels and filter by sample size (n). In addition, we perform conbination of five normalization methods and four zero replacement methods (for a total of 20 normalized datasets). In (2) Feature selection, we rank OTUs based on i) the number of times they are selected as important features by machine learning (ML) criteria, and ii) the greatest degree of difference on microbial networks reconstructed from samples of each class. We score OTUs based on whether they are selected by ML ($score=1$), by network comparison ($score=2$), both ($score=3$) or neither ($score=0$). In (3) Classification, the Venn diagram depicts the different types of predictors: microbiome (OTUs), environmental (Env), and the combination of both. The acronyms (e.g., All-OTU or OTU-S3+DS) correspond to different choices of predictors that are described in Table 2). Random forest and Bayesian neural network classification models are fitted on the different input predictors.
Results
@@
Overall performance of predictive models: Manual binarization causes inaccurate prediction of yield
@@
First, we identify the outcomes (disease or yield) that are accurately predicted across models (and are thus robust to prediction regardless of model choices), as well as the models that accurately predict across outcomes (and are thus the most powerful model alternatives). To do so, we aggregate the weighted F1 scores on data preprocessing choices such as normalization, zero replacement, and taxonomic levels for every model and every outcome. We employed RF and Bayesian NN models across various predictive scenarios. Figure 2 showcases the best performance of RF and Bayesian NN models. In the RF model, the integration of alpha diversity and soil chemistry data (referred to as Alpha+Soil) yields the most accurate predictions across all outcomes. Conversely, optimal performance for the Bayesian NN models is achieved by OTUs identified as significant by both ML and network comparison strategies (more details on feature selection are provided in Section ‘Effective preservation of predictive signal with different feature selection strategies’), alongside soil chemistry data (denoted as OTU-S3+Soil). These findings highlight the importance of integrating environmental information with microbiome data to enhance predictive power for disease outcomes. Further insights into the predictive capabilities of these models are presented in Figure S3 (RF) and Figure S4 (Bayesian NN), where a comprehensive analysis of all 14 models described in Table 2 is provided, offering valuable guidance for model selection and interpretation. Columns correspond to the six responses: four diseases and two yield outcomes. For a given panel (model in row and response in column), the boxplot corresponds to the different weighted F1 scores for every combination of normalizations/zero replacement strategies as well as different taxonomic levels (Table 3). For example, the boxplots for the ALL-OTU model (first row) include weighted F1 scores of the model fit on 20 normalization/zero replacement strategies, and 5 taxonomic levels (100 different weighted F1 scores per outcome). The dashed line in each panel corresponds to the average weighted F1 score of the model when fit with all random datasets (see Section ‘Prediction of potato disease from microbiome data’ and Figures S7 and S8 in Supplementary file). This line allows us to assess whether the real data has more predictive power than random data [1].
Figure 2: The most accurate predictions across all outcomes are achieved using alpha diversity and soil chemistry data (Alpha+Soil) for the RF model, whereas for the Bayesian NN models, optimal performance is observed when utilizing OTUs identified as important by both machine learning and network comparison strategies, in conjunction with soil chemistry data (OTU-S3+Soil). For a detailed presentation of all results, please refer to Supplementary Figures S3 (RF) and S4 (Bayesian NN).
@@
While these plots do not allow us to distinguish differences by taxonomic level or normalization/zero replacement strategy (more on that in the next subsections), we can identify outcomes (columns) that can be more accurately predicted across models (rows). Additionally, we can identify models that are capable of accurately predicting more outcomes. It is readily evident, for example, that the yield outcomes cannot be accurately predicted by any model as all the weighted F1 scores fall consistently below the dashed line. It is notable that the yield responses are those for which there is not a clear binarization strategy. Since we are artificially separating samples into the two classes (low and high yield) based on whether they are above or below the variety-specific median, samples on the boundary will in fact be very similar to each other, and thus, difficult to classify. Furthermore, the poor prediction of yield is not restricted to one data type (microbiome vs environmental) which also suggests that the prediction challenges arise from the binarization process rather than the model or set of predictors.
@@
Disease outcomes, on the contrary, display higher weighted F1 scores overall, and in particular, pitted scab displays weighted F1 scores that are consistently above the random prediction dashed line across different models. For the case of black scurf, even when the weighted F1 scores are very high, this is a deceiving result, as this disease outcome is highly imbalanced. This means that a naive model predicting all samples to belong to the majority class will have high prediction accuracy (see dashed line above 0.8 for random data). We investigate the prediction of black scurf more carefully with the augmented data that balances the proportion of both classes (Section ‘Robust prediction in imbalanced datasets with data augmentation’). Finally, given the poor performance on yield, we focus on the fine-grained description of results for the disease outcomes only for the remaining of the manuscript.
Prediction of potato disease from microbiome data
@@
No normalization or zero replacement strategy proven superior in terms of predictive power
@@
One of the goals of our study is to identify ideal data preprocessing steps that are guaranteed to maximize predictive power on the ML models. Our results, however, suggest that there is not a clear pattern of superiority among the normalization or zero replacement strategies, and that the selection of the data preprocessing steps needs to be data specific (and taxonomic level specific).
@@
Figure 3 shows the weighted F1 scores for pitted scab for different combinations of normalization and zero replacement strategies (x-axis) for the two types of models (RF and Bayesian NN). These analyses include all OTUs under the five taxonomic levels (different colors). Similar plots for other diseases are presented in Figures S9 to S13 in the Supplementary Material. As mentioned, we note that there is not an apparent pattern that points at any normalization or zero replacement strategy being superior to other alternatives across all taxonomic levels. In fact, there is considerable interaction between the normalization/zero replacement method and the taxonomic level. For example, for the RF model, the best result is achieved with Phylum level and cumulative sum scaling normalization with pseudo-zero replacement strategy (CSS+pseudo) or common sum scaling normalization without any zero replacement strategy (COM+none). Additionally, the rarefy+none and rarefy+multRepl strategies demonstrate good performance. For Bayesian NN, however, the best results are achieved with the common sum scaling normalization with the multiplicative zero replacement (COM+multRepl) for the Phylum level.
@@
For a given normalization/zero replacement strategy (x-axis), the variability in the scatterplot points indicates that taxonomic levels have an impact on the predictive power of the model. However, there is no discernible pattern indicating that a specific taxonomic level consistently provides more predictive signal. While the Phylum level appears to yield higher weighted F1 scores for certain normalization/zero replacement strategies, this trend is not observed across all strategies.
@@
When we compare the range of weighted F1 scores across normalization and zero replacement strategies, we see that the effect of the strategy is not negligible. For example, for a Phylum level, the lowest weighted F1 score is around 0.75 for centered log-ratio normalization with pseudo-zero replacement strategy (clr+pseudo) to around 0.9 for cumulative sum scaling normalization with pseudo-zero replacement strategy (CSS+pseudo). This implies that for a given taxonomic level, the resulted weighted F1 score will be highly influenced by the normalization and zero replacement strategy. Traditionally, microbiome researchers use the total sum scaling normalization without any zero replacement strategy (TSS+none) on their data which has a range of 0.80 to 0.90 weighted F1 scores for the RF model (0.8 to 0.85 for the Bayesian NN model) depending on the taxonomic level.
@@
The strong interaction effects of taxonomic level, normalization and zero replacement strategy prevent us from making recommendations about the best data preprocessing practices that can be generalizable to other datasets. We conclude by suggesting data practitioners to try different normalization and zero replacement strategies rather than just one, but see Section ‘Identifying general practices to predict potato disease from microbiome data using full model selection models’ for more recommendations.
@@
Figure 3: Weighted F1 scores (y-axis) for random forest and Bayesian neural network (Bayesian NN) models for the pitted scab disease under the 20 normalization/zero replacement strategies (x-axis). The lack of pattern prevents us from making recommendations of optimal strategies for microbiome data. We can conclude, however, that taxonomic levels, normalization and zero replacement strategies have an effect on the prediction accuracy of the models as evidenced by the broad range displayed by the points.
@@
Effective preservation of predictive signal with different feature selection strategies
@@
One of the standard steps in the ML pipeline is feature selection, especially for cases of high-dimensional data. We compare the ability to retain predictive signal of three feature selection strategies: standard importance score from ML methods, comparison of microbial network topologies and combination of both. More details on the feature selection strategies can be found in Methods. Figure 4 shows the weighted F1 scores for the two types of models (RF and Bayesian NN) on pitted scab disease (Scapbit) under different subsets of predictors: (1) all OTUs (ALL-OTU), (2) only OTUs that were identified as important by the ML strategy (OTU-S1), (3) only OTUs that were identified as important by the network comparison strategy (OTU-S2), (4) OTUs that were identified as important by both strategies (OTU-S3), or (5) OTUs that were not identified as important by neither strategy (OTU-S0). For fair comparison, we include the same number of predictors in OTU-S0 as in OTU-S3. Similar figures for other responses are shown in Figures S14 to S18 in the Supplementary Material.
@@
Again, we perceive strong interaction between taxonomic level and feature selection strategy. For the RF model, the highest weighted F1 score is achieved when including all OTUs (ALL-OTU) at Order level whereas for the Bayesian NN model, the highest weighted F1 score is achieved when including OTUs identified by the ML strategy (OTU-S1) at Genus level. RF models on all OTUs (ALL-OTU) have weighted F1 score above 0.8 in all taxonomic levels which suggest that this model could be a better alternative compared to Bayesian NN which is more computationally intensive. There are also smaller differences in RF models when comparing the performance on OTU-S3 (important OTUs) and ALL-OTU (all OTUs) which suggests that the feature selection strategy is sufficient to preserve the predictive signal in the data while reducing the number of predictors in the model. This is relevant for computationally intensive models such as Bayesian NN that do not allow the inclusion of all OTUs for certain taxonomic levels.
Robust prediction in imbalanced datasets with data augmentation
@@
To illustrate this, Figure 5 depicts the weighted F1 scores on original and augmented datasets for all yield and disease outcomes (x-axis) and both models (RF and Bayesian NN). The range of each box plot depicts the weighted F1 scores for 20 normalized datasets at each taxonomic level. We observe that black scurf and pitted scab can be reliably predicted across taxonomic levels as their median weighted F1 scores for all taxonomic orders is around $0.8$ when the models are fitted on the original datasets. As mentioned before, however, black scurf is highly imbalanced, so the results on the original data are not reliable. Fortunately, the median weighted F1 scores on augmented data (which is perfectly balanced by design) increases for both diseases, such that they are around $0.9$ for all taxonomic levels. These results suggest that data augmentation, especially in cases of highly imbalanced data, is an appropriate strategy that improve the robustness of the model, in some cases even increase the accuracy. One has to be careful, however, in that augmented data can yield certain models prohibited. For example, the Bayesian NN model could not be fit on the augmented datasets for Order, Family nor Genus levels due to computational limitations.
Figure 5: Boxplots with weighted F1 scores by model (rows): random forest (RF) and Bayesian neural network (Bayesian NN) for original and augmented OTU predictors (columns). The x-axis represents the yield and disease outcomes. The range of each box plot depicts the weighted F1 scores for 20 normalized datasets at each taxonomic level.
Identifying general practices to predict potato disease from microbiome data using full model selection models
@@
With a FMS [25, 26, 24] strategy, however, we are able to identify the choices that yield the highest measure of performance. More details on the FMS models can be found in Methods. Figures 6 and S19 show the FMS decision trees for the RF and Bayesian NN models on pitted scab disease, respectively. A FMS decision tree shows the different data preprocessing steps that yield different weighted F1 scores, so that practitioners can select the options that result in the highest predictive power. Here, we have five taxonomic levels, 20 normalization+zero replacement strategies, and 2 data augmentation options: no data augmentation (Aug=0) and data augmentation (Aug=1). Thus, in total, we have 200 data preprocessing options (20 normalization strategies times 5 taxonomic levels times 2 data augmentation).
@@
To interpret a FMS decision tree, each node corresponds to a specific step in the data preprocessing pipeline, for example, whether to perform data augmentation or not. If the condition is true, we follow the branch to the left; if the condition is false, we follow the branch to the right. At the top of the decision tree, we have the root which represents the data preprocessing step that has the greatest effect on model accuracy. At the bottom of the decision tree, we have the leaves with the average weighted F1 score of the model fitted on the data that satisfy all conditions towards the root. Each node also displays the percentage of data preprocessing options included in the node. For example, in Figure 6, the root node covers 100% of the options with average weighted F1 scores 0.865. The condition at the root node ($\text{Aug}=0$) represents the case of "no data augmentation". Thus, "true" (left of the root) means "no data augmentation", and "false" (right of the root) means "data augmentation". For simplicity, we denote the 20 normalization/zero replacement strategies as $\text{NM}_{i}$ for $i=1,\dots,20$. See Section ‘Data Filtering, Normalization, and Zero Replacement’ for a description on each normalization/zero replacement strategy.